Consider the following reaction:

 


PCl (g)+ Cl (g)    PCl  (g) K = 4.30 x 10-6

 

Initially, 2.50 M PCl3 and 1.40 M Cl2 are placed into a container and allowed to reach equilibrium. Determine the equilibrium concentrations of all the species


1
Expert's answer
2021-11-09T10:22:37-0500


4.30x10-6 = Kc





4.30x10-6 = [PCl5]/[PCl3]*[Cl2]


4.30x10-6 = [x]/[2.5]*[1.4]


x = 1.51x10-5